A cell of EMF E and internal resistance r supplies currents for the same time t through external resistance R1 and R2 respectively. if the heat produced in both cases is the same then the internal resistance is

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Current flowing in the first circuit isI_1=\dfrac{E}{R_1+r}

Power of heat dissipation in the first circuit=I_1^2R_1=\dfrac{E^2}{(R_1+r)^2}R_1

Since heat dissipated inttimeis same in both the circuits,

\dfrac{E^2}{(R_1+r)^2}R_1=\dfrac{E^2}{(R_2+r)^2}R_2
\implies r=\sqrt{R_1R_2}